OpenStudy (anonymous):

Which of the following is a factor of 8x3+ 1331? Which of the following is a factor of 8x3+ 1331? @Mathematics

OpenStudy (anonymous):

you have two cube roots therefore you can use the formula \[a^3-b^3=(a-b)(a^2+ab+b^2)\]

OpenStudy (anonymous):

if \[a^3=8x^3\] you need to find a therefore you're going to need to take the cube root of a^3 \[\sqrt[3]{8x^3} = \sqrt[3]{2x*2x*2x}\]

OpenStudy (anonymous):

so \[a=2x\] since you also need b you'll have to take the cube root of b^3 which will be \[\sqrt[3]{1331}=\sqrt[3]{11*11*11}=11\] b=11, so now all you need is to plug these into the equation above and that will be your answer \[a^3-b^3=8x^3-1331=(2x-11)((2x)^2+11(2x)+(11)^2)\]

OpenStudy (anonymous):

\[=(2x-11)(4x^2+22x+121)\]

OpenStudy (anonymous):

\[(2 x+11) \left(4 x^2-22 x+121\right)=8 x^3+1331 \]

OpenStudy (anonymous):

ah yes i had the signs flipped mine would be differences of cubes